Management of Mittel Rhein AG of Köln, Germany, would like to reduce the amount of time between when a customer places an order and when the order is shipped. For the first quarter of operations during the current year the following data were reported: Inspection time 0.7 days Wait time (from order to start of production) 16.6 days Process time 2.9 days Move time 0.8 days Queue time 4.3 days Required: 1. Compute the throughput time. (Round your answer to 1 decimal place.) 2. Compute the manufacturing cycle efficiency (MCE) for the quarter. (Round your percentage answer to nearest whole percent.) 3. What percentage of the throughput time was spent in non–value-added activities? (Round your percentage answer to nearest whole percent.) 4. Compute the delivery cycle time. (Round your intermediate calculations and final answer to 1 decimal place.) 5. If by using Lean Production all queue time during production is eliminated, what will be the new MCE? (Do not round intermediate calculations. Round your percentage answer to 1 decimal place.)
Answer #1

1. Throughput time (or Manufacturing cycle time) = Inspection Time + Process Time + Move Time + Queue Time = 0.7+2.9+0.8+4.3 = 8.7 days

Here Non-Value added time = Inspection Time + Move Time + Queue Time = 0.7+0.8+4.3 = 5.8 Days

Value added time = Process time = 2.9 days

2. Manufacturing cycle efficiency = Value Added time (Process time) / Manufacturing Cycle Time (Calculated in 1 above) = 2.9/8.7 = 33.33%

3. Percentage of non-value added activities in throughput time = 5.8 / 8.7 = 66.67%

4. Delivery Cycle time = Throughput time + Wait time = 8.7 days + 16.6 days = 25.3 Days

5. If queue time i.e. 4.3 days were eliminated, new Manufacturing cycle time = 0.7+2.9+0.8 = 4.4 days

Hence new MCE = 2.9/4.4 = 65.91%